Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60mph and train B is traveling at 80mph. Train A passes a station at 3:25pm, train B passes the same station at 3:40pm. At what time will Train B catch up to Train A
Answer by checkley77(12844)

D=60T FOR TRAIN a.
D=80(T-.25) FOR TRAIN B.
60T=80(T-.25)
60T=80T-20
60T-80T=-20
T=-20/-20
T=1 HOUR FOR TRAIN B TO CATCH TRAIN A.
PROOF:
60*1=80(1-.25
60=80*.75
60=60